3 questions:
one number is 3 more than another number, if two times the smaller plus 3 times the larger equals 29, what are the two numbers?
the length of a smaller table is 1 foot less than twice the width. the perimeter of this table is 16 feet. find the length and width of this table.
the length of a large table is 2 feet more than twice the width. the perimeter of the table is 22 feet. find the length and width of the table

Answer by jjthomson83(4)

Let the smaller number be x
Hence larger number is (x+3) since it's 3 more.
2x+3(x+3)=29
2x+3x+9=29
5x=20
x=4
x+3=4+3=7
Therefore the numbers are 4 and 7
2. Let the width of smaller table be x
Length of table=2x-1
Perimeter of table=2(x+2x-1)
2(3x-1)=16
3x-1=8
3x=9
x=3ft which is the width of table
Length is 2×3-1=5ft
3. 2(x+2x+2)=22
3x+2=11
x=3ft
Length=2×3+2=8ft